If 1 gallon of paint covers 400 square feet, about how many gallons of paint are needed to paint the outside of a spherical tank of diameter 20 feet?
Answer by stanbon(75887)

If 1 gallon of paint covers 400 square feet, about how many gallons of paint are needed to paint the outside of a spherical tank of diameter 20 feet?
-------------
Surface area of a sphere = 4(pi)r^2
Your Problem:
Surfact area = 4(pi)10^2 = 400pi sq ft.
--------------------------
400pi/400 = pi gallons of paint or approximately 3.14159 gallons.
===================
Cheers,
Stan H.